# Unraveling the Geometry of Visual Relational Reasoning

### v1

@@ Acknowledgments
Jiaqi Shang, Gabriel Kreiman, Haim Sompolinsky
Corresponding author. Email: hsompolinsky@mcb.harvard.edu
This PDF file includes:
Materials and Methods
Supplementary Text
Figures S1 to S8
Tables S1 to S2
Materials and Methods
SimplifiedRPM Dataset
The Raven’s Progressive Matrices (RPM) task  (?) is a widely used assessment of abstract relational reasoning, commonly applied in fields such as cognitive psychology and AI evaluation. Each RPM trial presents a 3×3 grid with the bottom-right panel missing. The subject must select the correct answer from eight options to complete the missing panel, ensuring that all three rows or columns in the matrix follow a consistent underlying relational rule.
Inspired by the RPM task, the SimplifiedRPM dataset was specifically designed to assess a model’s ability to generalize abstract relations across different object attributes (Fig. 1). Each row in the dataset follows a unique relational rule that determines how an object’s attribute values change across three panels (figure S6A). The dataset consists of $40$ relational rules, each derived by applying an abstract relation to a specific object attribute. These abstract relations fall into three categories: progression (constant, progression ±$\pm 2$, progression ±$\pm 1$), arithmetic (addition, subtraction), and logical (AND, OR, XOR). A detailed description of the abstract relations is provided in Table S1.
Each rule was applied to one of four object attributes: shape, size, shade, or number/position (Fig. S6). Number and position were grouped as a single attribute to avoid ambiguity. For example, a ”constant position” rule inherently fixes the object count, enforcing the constant number rule and conflicting with our requirement that each row in the task follows a unique relational rule. To ensure consistency with prior benchmarks, the dataset adopts the same object definitions and attribute specifications as the PGM dataset  (?). Shape is selected from seven geometric forms: circle, triangle, square, pentagon, hexagon, octagon, and star. The objects’ sizes are represented by their bounding boxes in the panel, with heights ranging from $20$ to $40$ pixels in ten evenly spaced increments. Shade is represented as a grayscale intensity, ranging from $[230,230,230]$ (close to white) to $[0,0,0]$ (black) in RGB values, with $10$ evenly spaced increments. The number attribute specifies the number of objects per panel, ranging from $1$ to $9$. Position refers to the spatial arrangement of objects. Each panel has a total size of $[160,160]$, and each object occupies a maximum bounding box of $[40,40]$. There are nine possible object locations, each defined by its centroid coordinates: $\{(20,20),(20,60),(20,100),(60,20),(60,60),(60,100),(100,20),(100,60),(100,100)\}$. Consequently, each row has dimensions $3\times 160\times 160$. When presented, the three panels are arranged horizontally.
Dataset generation followed a structured pipeline to ensure consistency and reproducibility. Each row was constructed to enforce a unique rule from our predefined set, eliminating ambiguity. To achieve this, we first sampled attribute values from valid options that conformed to the designated rule. Next, we randomly assigned values to the remaining attributes and resampled if a sampled value inadvertently conformed to an unintended rule. Once attribute values were finalized, they uniquely defined an input row. The corresponding row input was then generated by placing objects with the specified attributes in their designated positions. To maintain dataset balance, we generated 10,000 training rows, 1,000 validation rows, and 1,000 test rows per rule, yielding a total of 48,000 unique rows across all rules (40 rules $\times$ 1,200 samples). Example rows in the dataset are provided in figure S6B-D.
Neural Network Training
To evaluate a model’s ability to generalize abstract relations to unseen object attributes, we implement a rule-splitting strategy in which a subset of rules is withheld during training. Specifically, we randomly select five rules to be held out, ensuring that each abstract relation appears at most once in this set. The remaining $35$ rules are used for model training. After training, the model is evaluated based on its ability to infer the held-out rules. Fig. S6A provides an example of this rule split, with held-out rules highlighted in yellow and training rules in blue.
During training, models are optimized using supervised learning with a cross-entropy loss function. The models are trained on 10,000 examples per rule across $35$ training rules. Test error (Fig. 1C) is assessed based on classification accuracy using $1,000$ test examples per rule. These test examples differ from the $10,000$ training examples used during model training.
To evaluate model performance, we compute the Euclidean distance between the sample row and each of the two choice rows. A trial is considered correct if the correct choice row is closer to the sample row than the incorrect one. The model’s generalization error is then quantified as the proportion of incorrect trials averaged across all $500$ trials for each of the $20$ rule pairs (Fig. 1E).
We trained four architectures: ResNet-50, WReN, ViT, and SCL. ResNet-50  (?) is a convolutional neural network with residual connections. We employed the standard ResNet-50 architecture without modification. WReN  (?) explicitly models pairwise relations between objects and has been previously used for RPM tasks. We optimized the hyperparameters of its relation module and fully connected layers. ViT  (?) is a transformer-based model that operates on image patches as input tokens and captures global dependencies via self-attention. We split each of the three panels in a row into a $3\times 3$ grid, resulting in $27$ patches per row. We optimized key hyperparameters, including the number of layers, attention heads, and hidden dimensions. Finally, SCL  (?) was specifically designed for the RPM task, and we used the original architecture without modifications. Detailed model configurations, including layer-wise specifications and hyperparameter tuning, are provided in Supplementary Text.
Training was conducted using stochastic gradient descent with a learning rate of $0.1$, momentum of $0.9$, and weight decay of $5\times 10^{-4}$. To facilitate stable convergence, we applied a learning rate warmup for the first two epochs, followed by a plateau-based scheduler that reduced the learning rate by $0.2$ whenever the validation loss plateaued. Training continued until the classification errors across $1,000$ validation rows for each of the $35$ training rules stabilized. The final training durations for each model were: ResNet-50 ($100$ epochs), WReN ($60$ epochs), ViT ($200$ epochs), and SCL ($150$ epochs). All models were trained using a batch size of $512$.
Human Experiment
Twenty-five participants were recruited via Amazon Mechanical Turk (MTurk) to evaluate performance on the SimplifiedRPM dataset. To ensure data reliability, we set strict eligibility criteria, requiring an approval rating above $95\%$ and restricting participation to Masters-qualified workers. All participants provided informed consent, and the study protocol was approved by the Institutional Review Board at Children’s Hospital, Harvard Medical School. Compensation was performance-based, with participants earning more for higher accuracy.
Before the main experiment, participants read written instructions describing the task and completed three practice trials that introduced relational rules different from the $13$ rule pairs used in the main experiment. Participants received feedback and explanations of the underlying rules during these practice trials. In the main experiment, each participant completed 60 trials randomly selected from a larger pool of $130$ trials ($10$ trials per rule pair), encompassing $13$ distinct rule pairs of varying relational complexity (Table S2). Participants were shown images like those in Fig. 1A. The images remained on the screen until the participants made a choice. Participants responded by clicking on one of two choice rows and were allowed unlimited time to provide their answers. No performance feedback was given during the main experiment. Additional information, including illustrative example trials, can be found in figure S7.
Supplementary Text
Model Architecture and Hyperparameter Tuning
@@
We used the standard ResNet-50 architecture described by  (?) without any modifications. The architecture comprises residual blocks containing convolutional layers followed by batch normalization and ReLU activation. Specifically, the model begins with an initial convolutional layer, followed by four residual blocks, and concludes with a global average pooling layer and a fully connected layer for classification. To analyze the evolution of layer-wise geometry, we extracted features from the outputs of each residual block (Block 1–4) and the final output of the average pooling layer (avgpool).
@@
For hyperparameter tuning of WReN, we performed a grid search over the following parameters: (1) hidden dimension of the pairwise relational module $\{128,256,512\}$; (2) number of layers in the pairwise relational module $\{1,2,3,4\}$; (3) hidden dimension of the classification MLP $\{64,128,256\}$; and (4) number of layers in the classification MLP $\{1,2,3,4\}$. The optimal configuration, minimizing test error on three held-out rule splits, was found to be a pairwise relational module with $2$ layers of $512$ dimensions and a classification MLP with $3$ layers of $256$ dimensions.
@@
The Vision Transformer (ViT) architecture comprises multiple layers of multi-headed self-attention (MSA). In our implementation, we utilize a four-layer Transformer model, each containing $12$ attention heads and a hidden dimension of $96$. The model processes input row patches, embedding them into token representations with a learnable class token prepended to the sequence. These representations are then sequentially processed through self-attention layers to capture global contextual dependencies. For layer-wise geometric analysis, we extract the outputs from each attention layer (Attn1-4).
For hyperparameter tuning of ViT, we performed a grid search over the following parameters: (1) number of layers $\{4,8,12\}$; (2) number of attention heads $\{4,8,12,16\}$; (3) hidden dimension of the self-attention layers $\{48,96,192,384,768\}$. The optimal configuration, determined by minimizing test error, was found to be $4$ layers, each with $12$ heads and a hidden dimension of $96$.
@@
The Scattering Compositional Learner (SCL) model  (?) is a hierarchical architecture designed explicitly for RPM tasks (figure S8). Each input panel is first processed through a shared convolutional block, producing an individual panel embedding. These embeddings are then processed through three 1-D convolutional layers that operate across the three panels. Crucially, the same filter is applied consistently across the panel-wise feature dimensions, ensuring that abstract relations are detected in an attribute-invariant manner. Finally, the output is further refined using two MLP layers. For layer-wise geometric analysis, we extract representations from the convolutional block (CNN), the three 1-D convolutional layers (conv1-3), and the two MLP layers (MLP1–2).
For hyperparameter tuning of SCL, we performed a grid search over the following parameters: (1) the kernel size of the 1-D convolutional layers $\{1,2,4,8\}$; (2) the number of output channels, chosen from $\{16,32\}$ for the first layer, $\{32,64\}$ for the second layer, and $\{5,10\}$ for the final layer; (3) the hidden dimensions of the last two layers of the multi-layer perceptron (MLP) $\{400,1000\}$. The optimal configuration, determined by minimizing test error, was found to be a kernel size of $1$, output channel sizes of $\{64,32,5\}$, and an MLP hidden dimension of $400$.
Prototypical loss and SNR loss
For the SNR loss, we conduct a grid search over batch composition and learning rate. Following the setup used in cross-entropy training, we maintain a total batch size of $512$ and explore three batch compositions: $\{32\times 16,16\times 32,8\times 64\}$ ($m$ classes $\times$ $P$ samples per class). Additionally, we test three learning rates: $\{0.1,0.01,0.001\}$. Model performance is evaluated based on test error across three held-out rule splits, distinct from the $15$ used for training. The optimal configuration is $32\times 16$ with a learning rate of $0.01$.
@@
The prototypical loss  (?) is a distance-based loss function commonly used in few-shot learning. It generates rule prototypes, which act as representative representations for each rule. Like the SNR loss, each input batch consists of $P$ rows drawn from $m$ different relational rules. For each rule $r$, the $P$ rows are randomly split into two disjoint subsets: a support set $S_{r}$ and a query set $Q_{r}$. The prototype $z_{r}$ for rule $r$ is computed as the mean representation of its support set:
$z_{r}=\frac{1}{\mid S_{r}\mid}\sum_{x\in S_{r}}x$ (S1)
@@
$l_{\text{Prototypical}}=\sum_{r}\sum_{x\in Q_{c}}-\log\frac{\exp\left(-\parallel x-z_{c}\parallel^{2}\right)}{\sum_{c^{\prime}\neq c}\exp\left(-\parallel x-z_{c^{\prime}}\parallel^{2}\right)}$ (S2)
To optimize the Prototypical loss, we conducted a grid search over: (1) batch size $\{32\times 16,16\times 32,8\times 64\}$ ($m$ classes $\times$ $P$ samples per class); (2) learning rate $\{0.1,0.01,0.001\}$, and support set size $\{1,5\}$. Model performance was assessed using test accuracy on three held-out rule splits, distinct from the 15 used for training. The optimal configuration was found to be $32$ classes $\times$ $16$ samples per class, with a learning rate of $0.01$ and support set size $5$.
Figure S1: Model Test Errors Across Abstract Relation Types. (A-D) Test errors of four models on the $35$ training rules, evaluated across three abstract relation types: Progression (blue), Arithmetic (green), and Logical (red). The chance error is $0.97$. Each bar shows the mean classification error for rows of rules belonging to one of these relation types. Error bars represent the standard error of the mean (s.e.m.). Statistical significance was assessed using pairwise Wilcoxon tests, with asterisks indicating significance levels: *** ($p<0.001$).
@@
Figure S5: Comparison of Model Generalization Errors with Human Performance on the SimplifiedRPM Task. Changes in generalization signal (A-D), generalization dimensionality D (E-H), and the variance of held-out rule manifolds explained by the top 35 principal components of training rules (I-L) over training epochs. Results are averaged across 15 different held-out rule splits, with shaded areas representing the standard error of the mean (s.e.m.) for the four models (ResNet, WReN, ViT, and SCL). The red vertical line indicates the starting point of the plot in Fig. 6, marking the epoch at which the relational representation structure for the training rules stabilizes.
@@
Table S1: Relations in the SimplifiedRPM Dataset. The SimplifiedRPM dataset includes ten abstract relations organized into progression, arithmetic, and logical types. Each rule defines how attribute index values assigned to objects change across three panels. Each example consists of three lists, each representing a panel and specifying the assigned attribute values. For instance, $\left\{[1],[1],[1]\right\}$ represents a constant relation, with objects maintaining the attribute value one across all panels. Objects can also have multiple attribute values, as seen in the AND relation example $\left\{[1],[0,1],[1]\right\}$, where the second panel contains objects with attributes 0 and 1.
@@
# Correct Rule Incorrect Rule Human Error Response Time (sec)
1 Prog-1 Number AND Shape $0.11\pm 0.04$ $10.32\pm 1.33$
2 Prog+1 Number Const Number $0.14\pm 0.05$ $7.23\pm 1.13$
3 OR Shape Arith+ Number $0.15\pm 0.03$ $13.52\pm 1.08$
4 Prog+2 Number Const Number $0.18\pm 0.06$ $10.08\pm 1.45$
5 Arith- Number OR Shape $0.18\pm 0.06$ $13.19\pm 1.46$
6 XOR Shape AND Position $0.20\pm 0.04$ $18.57\pm 1.57$
7 Prog-2 Number XOR Position $0.22\pm 0.06$ $8.87\pm 1.23$
8 Arith+ Number OR Position $0.22\pm 0.06$ $12.13\pm 2.01$
9 Const Number Prog+2 Number $0.23\pm 0.04$ $6.08\pm 0.58$
10 AND Position Prog-2 Number $0.43\pm 0.07$ $14.95\pm 1.05$
11 AND Shape XOR Shape $0.47\pm 0.06$ $17.80\pm 1.07$
12 OR Position XOR Position $0.48\pm 0.07$ $20.62\pm 1.18$
13 XOR Position Arith- Number $0.51\pm 0.07$ $14.71\pm 1.02$

### v2

@@ Methods
Jiaqi Shang, Gabriel Kreiman, and Haim Sompolinsky
*Correspondence: hsompolinsky@mcb.harvard.edu
Appendix A Supplementary Methods
A.1 Model Architecture and Hyperparameter Tuning
@@
We used the standard ResNet-50 architecture described by [28] without any modifications. The architecture comprises residual blocks containing convolutional layers followed by batch normalization and ReLU activation. Specifically, the model begins with an initial convolutional layer, followed by four residual blocks, and concludes with a global average pooling layer and a fully connected layer for classification. To analyze the evolution of layer-wise geometry, we extracted features from the outputs of each residual block (Block 1–4) and the final output of the average pooling layer (avgpool).
@@
[…]
@@
The Vision Transformer (ViT) architecture comprises multiple layers of multi-headed self-attention (MSA). In our implementation, we utilize a four-layer Transformer model, each containing 121212 attention heads and a hidden dimension of 969696. The model processes input row patches, embedding them into token representations with a learnable class token prepended to the sequence. These representations are then sequentially processed through self-attention layers to capture global contextual dependencies. For layer-wise geometric analysis, we extract the outputs from each attention layer (Attn1-4).
[…]
@@
The Scattering Compositional Learner (SCL) model [9] is a hierarchical architecture designed explicitly for RPM tasks (figure S6). Each input panel is first processed through a shared convolutional block, producing an individual panel embedding. These embeddings are then processed through three 1-D convolutional layers that operate across the three panels. Crucially, the same filter is applied consistently across the panel-wise feature dimensions, ensuring that abstract relations are detected in an attribute-invariant manner. Finally, the output is further refined using two MLP layers. For layer-wise geometric analysis, we extract representations from the convolutional block (CNN), the three 1-D convolutional layers (conv1-3), and the two MLP layers (MLP1–2).
[…]
A.2 Prototypical loss and SNR loss
[…]
@@ A.2 Prototypical loss and SNR loss
The prototypical loss [33] is a distance-based loss function commonly used in few-shot learning. It generates rule prototypes, which act as representative representations for each rule. Like the SNR loss, each input batch consists of $P$ rows drawn from $m$ different relational rules. For each rule $r$, the $P$ rows are randomly split into two disjoint subsets: a support set $S_{r}$ and a query set $Q_{r}$. The prototype $z_{r}$ for rule $r$ is computed as the mean representation of its support set:
$z_{r}=\frac{1}{\mid S_{r}\mid}\sum_{x\in S_{r}}x$ (3)
@@
$l_{\text{Prototypical}}=\sum_{r}\sum_{x\in Q_{c}}-\log\frac{\exp\left(-\parallel x-z_{c}\parallel^{2}\right)}{\sum_{c^{\prime}\neq c}\exp\left(-\parallel x-z_{c^{\prime}}\parallel^{2}\right)}$ (4)
[…]
Appendix B Supplementary Figures
Figure S1: Model Test Errors Across Abstract Relation Types. (a-d) Test errors of four models on the 353535 training rules, evaluated across three abstract relation types: Progression (blue), Arithmetic (green), and Logical (red). The chance error is 0.970.970.97. Each bar shows the mean classification error for rows of rules belonging to one of these relation types. Error bars represent the standard error of the mean (s.e.m.). Statistical significance was assessed using pairwise Wilcoxon tests, with asterisks indicating significance levels: *** ($p<0.001$).
@@ Appendix B Supplementary Figures
Figure S8: Comparison of Model Generalization Errors with Human Performance on the SimplifiedRPM Task. Changes in generalization signal (a-d), generalization dimensionality D (e-h), and the variance of held-out rule manifolds explained by the top 35 principal components of training rules (i-l) over training epochs. Results are averaged across 15 different held-out rule splits, with shaded areas representing the standard error of the mean (s.e.m.) for the four models (ResNet, WReN, ViT, and SCL). The red vertical line marks the epoch at which the relational representation structure for the training rules stabilizes.
Appendix C Supplementary Tables
# Correct Rule Incorrect Rule Human Error Response Time (sec)
1 Prog-1 Number AND Shape 0.11±0.040.11$\pm$ 0.040.11 ± 0.04 10.32±1.3310.32$\pm$ 1.3310.32 ± 1.33
2 Prog+1 Number Const Number 0.14±0.050.14$\pm$ 0.050.14 ± 0.05 7.23±1.137.23$\pm$ 1.137.23 ± 1.13
3 OR Shape Arith+ Number 0.15±0.030.15$\pm$ 0.030.15 ± 0.03 13.52±1.0813.52$\pm$ 1.0813.52 ± 1.08
4 Prog+2 Number Const Number 0.18±0.060.18$\pm$ 0.060.18 ± 0.06 10.08±1.4510.08$\pm$ 1.4510.08 ± 1.45
5 Arith- Number OR Shape 0.18±0.060.18$\pm$ 0.060.18 ± 0.06 13.19±1.4613.19$\pm$ 1.4613.19 ± 1.46
6 XOR Shape AND Position 0.20±0.040.20$\pm$ 0.040.20 ± 0.04 18.57±1.5718.57$\pm$ 1.5718.57 ± 1.57
7 Prog-2 Number XOR Position 0.22±0.060.22$\pm$ 0.060.22 ± 0.06 8.87±1.238.87$\pm$ 1.238.87 ± 1.23
8 Arith+ Number OR Position 0.22±0.060.22$\pm$ 0.060.22 ± 0.06 12.13±2.0112.13$\pm$ 2.0112.13 ± 2.01
9 Const Number Prog+2 Number 0.23±0.040.23$\pm$ 0.040.23 ± 0.04 6.08±0.586.08$\pm$ 0.586.08 ± 0.58
10 AND Position Prog-2 Number 0.43±0.070.43$\pm$ 0.070.43 ± 0.07 14.95±1.0514.95$\pm$ 1.0514.95 ± 1.05
11 AND Shape XOR Shape 0.47±0.060.47$\pm$ 0.060.47 ± 0.06 17.80±1.0717.80$\pm$ 1.0717.80 ± 1.07
12 OR Position XOR Position 0.48±0.070.48$\pm$ 0.070.48 ± 0.07 20.62±1.1820.62$\pm$ 1.1820.62 ± 1.18
13 XOR Position Arith- Number 0.51±0.070.51$\pm$ 0.070.51 ± 0.07 14.71±1.0214.71$\pm$ 1.0214.71 ± 1.02
[…]